Online Survey Software x SV_egGCu5lIV7idaPb 13. Unemployment that arises as a result of the time it takes for unemployed people to locate a job utilizing their transferable skills is called O structural O cyclical O natural O frictional O full unemployment. 14. Look at the following data: The frictional unemployment rate is 2 percent, the structural unemployment rate is 3.5 percent, and the actual unemployment rate is 9.5 percent. The natural unemployment rate is percent. percent and the cyclical unemployment rate is O 1.5, 4 O 55,4 O 1.5, 6 O 6, 15 15. At the time of Kelseys 20 year high school reunion she was making $50,000 and the CPI was 120. Now that it is time for her to attend her 30 year high school reunion, Kelseys income has risen to $97,000 and the CPl is 230. At her 30 year reunion, can Kelsey rightfully brag that her real income has risen since the last time she saw her former classmates ten years ago? O Yes, Kelseys real income rose during that 10 year period O No, Kelseys real income fell during that 10 year period O No, Kelseys real income remained constant during that 10 year period O It is impossible to determine what happened to Kelseys real income.
Answer #1

13. Ans: Frictional

Explanation:

Frictional unemployment occurs in any economy because people being in the process of moving from one job to another job.

14. Ans: 5.5 ; 4

Explanation:

Natural unemployment rate = 2% + 3.5% = 5.5%

Cyclical unemployment rate = 9.5% - 5.5% = 4%

15. Ans: Yes, Kelsey's real income rose during that 10 years period.

Explanation:

Intial Real income = (50,000 / 120) * 100 = $41,666.67

Real income after 10 years = (97,000 / 230) * 100 = 42,174